Solve the following systems of equations using any method.
( c ) 2.4 x = 2.2 y – 3.2
1.6 x + 1.2 y = 3.2

Answer by mananth(16946)

2.4 x = 2.2 y – 3.2
1.6 x + 1.2 y = 3.2
whenever there are decimal equations
multiply the equations by 10 if there is 1 decimal and by 100 if there are 2 decimals. It makes your life easy.
..
There is only one decimal
24x-22y=-32............. equation 1
16x+12y=32 ............. equation 2
..
multiply equation 1 by 2 and equation 2 by 3 and subtract to eliminate x
48x-44y=-64
48x+36y=96
..
48x-44y-48x-36y=-64-96
-80x=-160
x=2
plug the value of x in equation 1
48-22y=-32
-22y=-80
y=3.64